OpenStudy (anonymous):

3/4 - 1/2 (1/3)+(3/7)=

OpenStudy (anonymous):

I came up with 43/84? can anyone tell me if I'm correct? lottsa thx

OpenStudy (anonymous):

i came up with 85/84 for some reason

OpenStudy (gw2011):

91rt is correct. Here is how the answer was arrived at: 3/4-1/6+3/7=63/84-14/84+36/84=85/84

OpenStudy (anonymous):

thank you guys!!!

OpenStudy (gw2011):

You're welcome